# TDHook: A Lightweight Framework for Interpretability

### v1

@@ Abstract
Interpretability of Deep Neural Networks (DNNs) is a growing field driven by the study of vision and language models. Yet, some use cases, like image captioning, or domains like Deep Reinforcement Learning (DRL), require complex modelling, with multiple inputs and outputs or use composable and separated networks. As a consequence, they rarely fit natively into the API of popular interpretability frameworks. We thus present TDHook, an open-source, lightweight, generic interpretability framework based on tensordict and applicable to any torch model. It focuses on handling complex composed models which can be trained for Computer Vision, Natural Language Processing, Reinforcement Learning or any other domain. This library features ready-to-use methods for attribution, probing and a flexible get-set API for interventions, and is aiming to bridge the gap between these method classes to make modern interpretability pipelines more accessible. TDHook is designed with minimal dependencies, requiring roughly half as much disk space as transformer_lens, and, in our controlled benchmark, achieves up to a $\times$2 speed-up over captum when running integrated gradients for multi-target pipelines on both CPU and GPU. In addition, to value our work, we showcase concrete use cases of our library with composed interpretability pipelines in Computer Vision (CV) and Natural Language Processing (NLP), as well as with complex models in DRL.
@@ 1 Introduction
The increasing complexity of Deep Neural Networks (DNNs) has raised significant challenges in understanding their decision-making processes and internal representations. Advancements in large language models (Radford & Narasimhan, 2018), computer vision systems (Simonyan & Zisserman, 2014), and other deep learning architectures have further exacerbated the interpretability challenge by mixing modalities and domains, making it increasingly difficult for researchers and practitioners to debug, let alone trust, these systems. While the field of explainable AI has grown substantially in recent years, existing efforts in model interpretability predominantly focus on specific domains and architectures. Methods for attribution (Simonyan et al., 2013; Sundararajan et al., 2017), latent manipulation (Zou et al., 2023; Cunningham et al., 2023), and weights-based analysis (Dunefsky et al., 2024) have been well-developed, and frameworks like captum (Kokhlikyan et al., 2020), zennit (Anders et al., 2021), and pyvene (Wu et al., 2024) provide specialized tools for these techniques.
@@
  • Introduction of a new interpretability framework, TDHook [^1] The library is available at https://github.com/Xmaster6y/tdhook, based on torch (Ansel et al., 2024) and tensordict (Bou et al., 2023).
@@
  • Concrete use cases of the TDHook in various domains with extension of multi-step interpretability pipelines.
@@
Figure 1: Schematic view of the TDHook framework architecture. The target to interpret is a torch module or a TensorDictModule wrapped in a HookedModel. This object exposes different APIs, like the get-set API described in Section 3.3, while its forward can be modified to fit the needs of different interpretability techniques, see an example in Section 3.2. The artifacts produced and consumed by the models and methods are materialised as TensorDict objects. More details about the design principles are given in Section 3.1, while interactions between the main components of the framework are given in Appendix C.
@@ 2.1 Interpreting Deep Neural Networks
As we are interested in methods that can be directly applied to a Deep Neural Network (DNN) model, we take a post-hoc approach. In Computer Vision (CV) and Natural Language Processing (NLP), many approaches were developed in order to study pre-trained models (Simonyan & Zisserman, 2014; He et al., 2015; Radford & Narasimhan, 2018). In order to improve code modularity in the TDHook library, we use a categorisation of these methods based on their data stream and produced artifacts.
@@
These are typical methods used in CV to understand convolutional networks by visualising important pixels, i.e. saliency maps, (Simonyan et al., 2013). The key idea is to compute the sensitivity of the output or any intermediate representation with respect to the input or other intermediate representation. The first developed class of methods were gradient-based (Simonyan et al., 2013), with derived methods like integrated gradients (Sundararajan et al., 2017) or grad-CAM (Selvaraju et al., 2016), that are mathematically grounded with sensitivity. Certain methods compute importance maps by occluding part of the input (Zeiler & Fergus, 2013) or, more generally, perturbing the input (Covert et al., 2020). Other backpropagation methods like LRP locally define relevance rules, using deep-Taylor expansions, which enable global attribution (Montavon et al., 2015; Bach et al., 2015). Recent works in NLP focus on the Transformer architecture and its attention mechanism (Vaswani et al., 2017), providing token-level insights (Wiegreffe & Pinter, 2019; Achtibat et al., 2024). Finally, some methods go beyond single prediction explanation and aim to provide more conceptual explanations, like the computation of pre-images from activation maximisation (Mahendran & Vedaldi, 2015), or the finding of maximally activating samples in a dataset (Chen et al., 2020).
@@
This class of methods techniques extends the interpretability of concepts and features by exploring the internal representations learned by models. Latent concepts were introduced in CV with (Kim et al., 2018), where concepts are defined as latent space directions, typically computed with linear probes (Alain & Bengio, 2018; Dreyer et al., 2023b). This field extends the notion of prototypes like perturbed images (Ribeiro et al., 2018), cropped images (Dreyer et al., 2023a) or pre-images (Mahendran & Vedaldi, 2015). These methods were later expanded by the field of representation engineering (Zou et al., 2023), where such latent features enable locating, editing, erasing or decoding models’ knowledge (Meng et al., 2022; Belrose et al., 2023; Ghandeharioun et al., 2024). Other works enable causally modifying or analysing the model outputs (Rimsky et al., 2023; Kram’ar et al., 2024), or use sparse autoencoders to elicit interpretable features in language models (Cunningham et al., 2023).
@@
More focused on data-independent interpretability, these methods provide a more granular understanding of model internals by examining pathways and dependencies between models’ components, usually neurons or attention heads, and aim to remove the data dependency from explanations in order to provide out-of-distribution explanations. Viewing models’ components as circuits was first proposed for CNNs (Olah et al., 2020) before being formalised for Transformers (Elhage et al., 2021). These circuits revealed peculiar models’ components that learned precise mechanisms like induction (Olsson et al., 2022). Using specific datasets, relevant circuits can be automatically discovered (Conmy et al., 2023), and can also involve larger models’ components at the layer scale (Dunefsky et al., 2024). In this framework, operations on the weights become natural, like editing using task vectors (Ilharco et al., 2022; Hendel et al., 2023). As a consequence, it is also possible to use explanations, from attribution or circuit discovery, to selectively prune models (Yeom et al., 2019; Pochinkov & Schoots, 2024).
@@ 2.2 Interpretability Frameworks
A range of open-source libraries has emerged to operationalise some of the previously described techniques, each with its own focus and design philosophy. captum (Kokhlikyan et al., 2020) is the most widely-used torch interpretability library, in terms of downloads and GitHub stars, and concentrates on attribution, providing more than thirty gradient and perturbation-based explainers. zennit (Anders et al., 2021), the de facto library for Layer-wise Relevance Propagation methods, offers a unified interface for relevance-based explanations. In the causal-analysis corner, pyvene (Wu et al., 2024) provides configuration-based composable interventions while nnsight (Fiotto-Kaufman et al., 2024) exposes a general intervention API, based on a getters and setters (get-set API), for altering models’ computations at inference time. inseq (Sarti et al., 2023) specialises in sequence-generation models attribution, most notably by porting the methods from captum. Finally, transformer_lens (Nanda & Bloom, 2022) targets Transformer models and eases their study by providing a unified access to different model architectures.
@@
Table 1: Comparison of interpretability frameworks on selected features. Any Torch: compatibility with arbitrary PyTorch models; TensorDict: support for TensorDict and TensorDictModule; Get-Set: availability of a generic intervention API; Composable: ability to compose different methods together; Methods: number of ready-to-use interpretability methods; Coverage %: percentage of the codebase covered by tests.
@@ 3.1 Design Principles
Similarly to captum (Kokhlikyan et al., 2020), we focus our efforts on making interpretability methods easy to use, and provide a wide range of methods, ready to be used even by non-expert users. This strength is especially important in order to democratise the interpretability of DDNs. In that respect, TDHook is not only a generic interpretability framework but also a pragmatic and easy way to integrate interpretability methods. An example using a ready-to-use method is given in Figure 2 and a list of the currently implemented methods is given in Appendix C.
@@ 3.2 Ready-to-use Methods
Ready-to-use methods are a powerful adjunct to any library as they enable non-experts to apply interpretability methods. These methods encapsulate common interpretability workflows into simple function calls, abstracting away the complexity of manually setting up hooks, interventions, and data processing pipelines. By providing pre-configured implementations for methods like saliency mapping, latent manipulation, and weights manipulation, users can focus on their research questions rather than implementation details. The underlying complexity of hooking mechanisms, tensor manipulation, and result formatting is handled automatically, making interpretability accessible to researchers with varying levels of technical expertise. This is especially relevant for cross-disciplinary research where DNNs, used as black-box tools, would benefit from being explained. Figure 2 illustrates the execution flow for a saliency method, showing how the context factory prepares the model and how attributions are retrieved. For comprehensive implementation details and code examples, we refer the reader to Appendix C and the supplementary materials.
@@ 3.3 General Get-Set API
As already proposed by (Fiotto-Kaufman et al., 2024), the get-set API is a versatile way to interact with a model. This pattern separates the definition of interventions (with getters and setters) from their execution and result retrieval. We adopt a more explicit approach using HookedModel methods that provide fine-grained control over model interventions and state retrieval. The get-set API allows researchers to first define their intervention scheme, specifying which layers to hook, what modifications to apply, and which intermediate states to capture, before executing the model. This separation enables complex intervention strategies to be built incrementally and ensures that all interventions are applied consistently during a single forward pass and a potential backward pass. The API returns cache proxies that act as placeholders for future values, which are then populated during execution and can be resolved afterwards to retrieve the actual intermediate representations. The model can also be executed in a run context offering the same API, triggering the registered hooks and populating the cache with the intermediate representations. This context automatically handles hooks deregistration, temporary cache configuration and offers other capabilities like stopping the computation at any layer. Figure 3 demonstrates the intervention workflow, showing how hooks are registered during the setup phase and how intermediate states are retrieved after execution. Implementation details can be found in Appendix C and in the supplementary materials.
@@ 4 Benchmarking
In order to fairly compare the different libraries, we use the tasks which are showcased as a "Get Started" task in their respective documentation, and we reproduce them in TDHook, checking the consistency of the results [^2] Caching in transformer_lens is not reproduced exactly due to the model modifications needed.. For each task, we run multiple experiments with different seeds, model sizes, batch sizes and other parameters, consisting of more than a hundred variations. For reproducibility purposes, we give more details in Appendix A and provide the complete code in the supplementary material. We also depict the exact hardware configuration used for the benchmark in Appendix A. In Figure 4, we report the aggregated results of our benchmark. We measure the relative performance of each library against TDHook across different metrics (time, RAM, VRAM) and different setups (CPU only or CPU+GPU). Each task run is performed in isolation and always starts with a "spawn" (model setup or other preparations without computation) for which the metrics are also reported. This kind of benchmark is useful to stress-test the different libraries and to highlight the different trade-offs between them. For example, while Captum is optimised in terms of memory usage, it is slower due to its gradient computation strategy [^3] Captum only computes the gradient with respect to scalar values instead of vector values., it does not scale well with the batch size. TransformerLens has a setup overhead in terms of memory usage in order to convert the model to a unified format; it does not scale well with the model size.
@@ 5.1 Complex Pipelines
Contrary to traditional attribution methods that focus on explaining the output of the model, e.g. the class predicted, concept attribution methods focus on explaining the output with respect to specific concepts, illustrated in Figure 6. Most of these methods rely on conditionally propagating the attribution with respect to specific channels, like conductance for integrated gradients (Dhamdhere et al., 2018) or concept relevance propagation for LRP (Achtibat et al., 2022). These channel concepts are usually discovered or visualised using activation maximisation (Chen et al., 2020) or relevance maximisation (Achtibat et al., 2022). But since concept activation vectors (CAVs) generalise these concepts, we introduce a propagation method based on the TCAV scores (Kim et al., 2018). Different methods for producing CAVs were already compared in (Dreyer et al., 2023b), but not in the context of concept attribution.
@@
Concept attribution is complex, as it is a composed method, and it is important to be able to compare across a variety of possible combinations of methods. As all methods are implemented in TDHook, it is possible to easily compare them by simply switching the method used. In order to evaluate the methods, we randomise the input according to the relevance scores, i.e. pixel flipping (Bach et al., 2015), and record the change in the class logit and the concept probability. The results, reported in Figure 7, show that using the TCAV scores gives attribution that is more faithful to the concept. Yet, qualitative analysis suggests that the obtained attribution is not localised and other metrics should be used to evaluate the quality of the attribution, like localisation or robustness metrics (Hedström et al., 2022).
@@
Attribution patching, also known as causal mediation analysis (Vig et al., 2020), is a method to understand the contribution of different components of the model to the output. It provides a powerful causal analysis but is costly to compute, as it requires as many forward passes as there are components to patch. Attribution patching (Nanda, 2022) was proposed as an alternative using a first-order approximation, only requiring two forward passes and one backward pass. Recent work has replaced the gradient by a relevance score obtained using LRP, i.e. RelP (Jafari et al., 2025). We explore the generalisability of this idea, and the results of the RelP method are sensitive to the LRP rules used, notably the results using the AH-rule presented in the paper, but not used for the experiments, are presented in Appendix B. The results in Figure 8 show that only methods closely related to the gradient are relevant for attribution patching. In fact, the rules used by RelP correct local gradient shatters, but do not conserve the relevance throughout propagation like traditional LRP rules.
Figure 8: Pearson correlation coefficients between different attribution patching methods and the activation patching method. This metric does not account for amplitude differences between the methods.

### v2

@@ Abstract
Interpretability of Deep Neural Networks (DNNs) is a growing field driven by the study of vision and language models. Yet, some use cases, like image captioning, or domains like Deep Reinforcement Learning (DRL), require complex modelling, with multiple inputs and outputs or use composable and separated networks. As a consequence, they rarely fit natively into the API of popular interpretability frameworks. We thus present TDHook, an open-source lightweight, generic interpretability framework based on tensordict and applicable to any torch model. It focuses on handling complex composed models which can be trained for Computer Vision (CV), Natural Language Processing (NLP), Reinforcement Learning or any other domain. This library features ready-to-use methods for attribution, probing and a flexible get-set API for interventions, and is aiming to bridge the gap between these method classes to make modern interpretability pipelines more accessible. TDHook is designed with minimal dependencies, requiring roughly half as much disk space as transformer_lens, and, in our controlled benchmark, achieves at least a $\times$2 speed-up over captum when running integrated gradients for multi-target pipelines on both CPU and GPU. In addition, to value our work, we showcase concrete use cases of our library with composed interpretability pipelines in CV and NLP, as well as with complex models in DRL.
@@ 1 Introduction
The increasing complexity of Deep Neural Networks (DNNs) has raised significant challenges in understanding their decision-making processes and internal representations. Advancements in large language models (Radford and Narasimhan, 2018), computer vision systems (Simonyan and Zisserman, 2014), and other deep learning architectures have further exacerbated the interpretability challenge by mixing modalities and domains, making it increasingly difficult for researchers and practitioners to debug, let alone trust, these systems. While the field of explainable AI has grown substantially in recent years, existing efforts in model interpretability predominantly focus on specific domains and architectures. Methods for attribution (Simonyan et al., 2013; Sundararajan et al., 2017), latent manipulation (Zou et al., 2023; Cunningham et al., 2023), and weights-based analysis (Dunefsky et al., 2024) have been well-developed, and frameworks like captum (Kokhlikyan et al., 2020), zennit (Anders et al., 2021), and pyvene (Wu et al., 2024) provide specialised tools for these techniques.
@@
  • Introduction of a new interpretability framework, TDHook [^1] The library is provided as supplementary material, based on torch (Ansel et al., 2024) and tensordict (Bou et al., 2023).
@@
  • Concrete use cases of TDHook applied to language, vision, and DRL models, illustrating multi-step interpretability pipelines.
@@
Figure 1: Schematic view of the TDHook framework architecture. The target to interpret is a torch module or a TensorDictModule wrapped in a HookedModel. This object exposes different APIs, like the get-set API described in Section 3.3, while its forward can be modified to fit the needs of different interpretability techniques, see an example in Section 3.2. The artifacts produced and consumed by the models and methods are materialised as TensorDict objects. More details about the design principles are given in Section 3.1, while interactions between the main components of the framework are given in Appendix D.
@@ 2.1 Interpreting Deep Neural Networks
As we are interested in methods that can be directly applied to a Deep Neural Network (DNN) model, we take a post-hoc approach. In Computer Vision (CV) and Natural Language Processing (NLP), many approaches were developed in order to study pre-trained models (Simonyan and Zisserman, 2014; He et al., 2015; Radford and Narasimhan, 2018). In order to improve code modularity in the TDHook library, we use a categorisation of these methods based on their data stream and produced artifacts.
@@
These are typical methods used in CV to understand convolutional networks by visualising important pixels, i.e. saliency maps, (Simonyan et al., 2013). The key idea is to compute the sensitivity of the output or any intermediate representation with respect to the input or other intermediate representation. The first developed class of methods were gradient-based (Simonyan et al., 2013), with derived methods like integrated gradients (Sundararajan et al., 2017) or grad-CAM (Selvaraju et al., 2016), that are mathematically grounded with sensitivity. Certain methods compute importance maps by occluding part of the input (Zeiler and Fergus, 2013) or, more generally, perturbing the input (Covert et al., 2020). Other backpropagation methods like LRP locally define relevance rules, using deep-Taylor expansions, which enable global attribution (Montavon et al., 2015; Bach et al., 2015). Recent works in NLP focus on the Transformer architecture and its attention mechanism (Vaswani et al., 2017), providing token-level insights (Wiegreffe and Pinter, 2019; Achtibat et al., 2024). Finally, some methods go beyond single prediction explanation and aim to provide more conceptual explanations, like the computation of pre-images from activation maximisation (Mahendran and Vedaldi, 2015), or the finding of maximally activating samples in a dataset (Chen et al., 2020).
@@
This class of methods techniques extends the interpretability of concepts and features by exploring the internal representations learned by models. Latent concepts were introduced in CV with (Kim et al., 2018), where concepts are defined as latent space directions, typically computed with linear probes (Alain and Bengio, 2018; Dreyer et al., 2023b). This field extends the notion of prototypes like perturbed images (Ribeiro et al., 2018), cropped images (Dreyer et al., 2023a) or pre-images (Mahendran and Vedaldi, 2015). These methods were later expanded by the field of representation engineering (Zou et al., 2023), where such latent features enable locating, editing, erasing or decoding models’ knowledge (Meng et al., 2022; Belrose et al., 2023; Ghandeharioun et al., 2024). Other works enable causally modifying or analysing the model outputs (Rimsky et al., 2023; Kram’ar et al., 2024), or use sparse autoencoders to elicit interpretable features in language models (Cunningham et al., 2023).
@@
These methods focus on interventions and analysis of model parameters themselves, rather than activations or gradients (Saphra and Wiegreffe, 2024). They provide a more granular understanding of model internals by examining pathways and dependencies between models’ components, usually neurons or attention heads, and aim to remove the data dependency from explanations in order to provide out-of-distribution explanations. Viewing models’ components as circuits was first proposed for CNNs (Olah et al., 2020) before being formalised for Transformers (Elhage et al., 2021). These circuits revealed peculiar models’ components that learned precise mechanisms like induction (Olsson et al., 2022). Using specific datasets, relevant circuits can be automatically discovered (Conmy et al., 2023), and can also involve larger models’ components at the layer scale (Dunefsky et al., 2024). In this framework, operations on the weights become natural, like editing using task vectors (Ilharco et al., 2022; Hendel et al., 2023). As a consequence, it is also possible to use explanations, from attribution or circuit discovery, to selectively prune models (Yeom et al., 2019; Pochinkov and Schoots, 2024).
@@ 2.2 Interpretability Frameworks
A range of open-source libraries has emerged to operationalise some of the previously described techniques, each with its own focus and design philosophy. captum (Kokhlikyan et al., 2020) is the most widely-used torch interpretability library, in terms of downloads and GitHub stars, and concentrates on attribution, providing more than thirty gradient and perturbation-based explainers. zennit (Anders et al., 2021), the de facto library for Layer-wise Relevance Propagation methods, offers a unified interface for relevance-based explanations. In the causal-analysis corner, pyvene (Wu et al., 2024) provides configuration-based composable interventions while nnsight (Fiotto-Kaufman et al., 2024) exposes a general intervention API, based on a getters and setters (get-set API), for altering models’ computations at inference time. inseq (Sarti et al., 2023) specialises in sequence-generation models attribution, most notably by porting the methods from captum. Finally, transformer_lens (Nanda and Bloom, 2022) targets Transformer models and eases their study by providing a unified access to different model architectures.
@@
Table 1: Comparison of interpretability frameworks on selected features. Any Torch: compatibility with arbitrary PyTorch models; TensorDict: support for TensorDict and TensorDictModule; Get-Set: availability of a generic intervention API; Composable: ability to compose different methods together; Methods: number of ready-to-use interpretability methods; Coverage %: percentage of the codebase covered by automated tests.
@@ 3.1 Design Principles
Similarly to captum (Kokhlikyan et al., 2020), we focus our efforts on making interpretability methods easy to use, and provide a wide range of methods, ready to be used even by non-expert users. This strength is especially important in order to democratise the interpretability of DDNs. In that respect, TDHook is not only a generic interpretability framework but also a pragmatic and easy way to integrate interpretability methods. An example using a ready-to-use method is given in Figure 2 and a list of the currently implemented methods is given in Appendix D.
@@ 3.2 Ready-to-use Methods
Ready-to-use methods are a powerful adjunct to any library as they enable non-experts to apply interpretability methods. These methods encapsulate common interpretability workflows into simple function calls, abstracting away the complexity of manually setting up hooks, interventions, and data processing pipelines. By providing pre-configured implementations for methods like saliency mapping, latent manipulation, and weights manipulation, users can focus on their research questions rather than implementation details. The underlying complexity of hooking mechanisms, tensor manipulation, and result formatting is handled automatically, making interpretability accessible to researchers with varying levels of technical expertise. This is especially relevant for cross-disciplinary research where DNNs, used as black-box tools, would benefit from being explained. Figure 2 illustrates the execution flow for a saliency method, showing how the context factory prepares the model and how attributions are retrieved. For comprehensive implementation details and code examples, we refer the reader to Appendix D and the supplementary materials.
@@ 3.3 General Get-Set API
As already proposed by (Fiotto-Kaufman et al., 2024), the get-set API is a versatile way to interact with a model. This pattern separates the definition of interventions (with getters and setters) from their execution and result retrieval. We adopt a more explicit approach using HookedModel methods that provide fine-grained control over model interventions and state retrieval. The get-set API allows researchers to first define their intervention scheme, specifying which layers to hook, what modifications to apply, and which intermediate states to capture, before executing the model. This separation enables complex intervention strategies to be built incrementally and ensures that all interventions are applied consistently during a single forward pass and a potential backward pass. The API returns cache proxies that act as placeholders for future values, which are then populated during execution and can be resolved afterwards to retrieve the actual intermediate representations. The model can also be executed in a run context offering the same API, triggering the registered hooks and populating the cache with the intermediate representations. This context automatically handles hooks deregistration, temporary cache configuration and offers other capabilities like stopping the computation at any layer. Figure 3 demonstrates the intervention workflow, showing how hooks are registered during the setup phase and how intermediate states are retrieved after execution. Implementation details can be found in Appendix D and in the supplementary materials.
@@ 4 Benchmarking
In order to fairly compare the different libraries, we use the tasks which are showcased as a "Get Started" task in their respective documentation, and we reproduce them in TDHook, checking the consistency of the results [^2] Caching in transformer_lens is not reproduced exactly due to the model modifications needed.. For each task, we run multiple experiments with different seeds, model sizes, batch sizes and other parameters, consisting of more than a hundred variations. Our experiments include models ranging from small MLPs to transformers (using the GPT-2 family), with up to 1B parameters. The experiments were conducted with batch sizes up to 100k for MLPs, in order to explore the limits of common use cases like activation caching. For reproducibility purposes, we give more details in Appendix A and provide the complete code in the supplementary material. We also depict the exact hardware configuration used for the benchmark in Appendix A. In Figure 4, we report the aggregated results of our benchmark. We measure the relative performance of each library against TDHook across different metrics (time, RAM, VRAM) and different setups (CPU only or CPU+GPU). Each task run is performed in isolation and always starts with a "spawn" (model setup or other preparations without computation) for which the metrics are also reported. This kind of benchmark is useful to stress-test the different libraries and to highlight the different trade-offs between them in terms of resource consumption. For example, while Captum is optimised in terms of memory usage, it is slower due to its gradient computation strategy [^3] Captum only computes the gradient with respect to scalar values instead of vector values., it does not scale well with the batch size. TransformerLens has a setup overhead in terms of memory usage in order to convert the model to a unified format; it does not scale well with the model size.
@@ 5.1 Complex Pipelines
Contrary to traditional attribution methods that focus on explaining the output of the model, e.g. the class predicted, with respect to the input, concept attribution methods focus on explaining the output with respect to specific concepts, illustrated in Figure 6. Most of these methods rely on conditionally propagating the attribution with respect to specific channels, like conductance for integrated gradients (Dhamdhere et al., 2018) or concept relevance propagation for LRP (Achtibat et al., 2022). These channel concepts are usually discovered or visualised using activation maximisation (Chen et al., 2020) or relevance maximisation (Achtibat et al., 2022). But since concept activation vectors (CAVs) generalise these methods, we introduce a propagation method based on the TCAV scores (Kim et al., 2018). Different methods for producing CAVs were already compared in (Dreyer et al., 2023b), but not in the context of concept attribution.
@@
Concept attribution is complex, as it is a composed method, and it is important to be able to compare across a variety of possible combinations of methods. As all methods are implemented in TDHook, it is possible to easily compare them by simply switching the method used, enabling users to chose their prefered notion of concept or attribution. For our experiments, we use the VGG16 model (Simonyan and Zisserman, 2014), and extracted concepts like "striped" using the texture dataset from (Cimpoi et al., 2013). In order to evaluate the methods, we randomise the input according to the relevance scores, i.e. pixel flipping (Bach et al., 2015), and record the change in the class logit and the concept probability. The results, reported in Figure 7, show that using the TCAV scores gives attribution that is more faithful to the concept. TCAV marginally outperforms the other methods according to pixel-flipping evaluation, as shown in Figure 7. Yet, qualitative analysis suggests that the obtained attribution is not localised, e.g., showing relevance spread beyond the lemur’s tail. Therefore other metrics should be used to evaluate the quality of the attribution, like localisation or robustness metrics (Hedström et al., 2022).
@@
Activation patching, also known as causal mediation analysis (Vig et al., 2020), is a method to understand the contribution of different components of the model to the output. It provides a powerful causal analysis but is costly to compute, as it requires as many forward passes as there are components to patch. Attribution patching (Nanda, 2022) was proposed as an alternative using a first-order approximation, only requiring two forward passes and one backward pass. Recent work has replaced the gradient by a relevance score obtained using LRP, i.e. RelP (Jafari et al., 2025). We explore the generalisability of this idea by replacing gradients by latent relevances. And show that the results of the RelP method are sensitive to the LRP rules used, notably the results using the AH-rule, introduced in the paper Jafari2025RelPFA but not used for the experiments, are presented in Appendix B. The results in Figure 8 show that only methods closely related to the gradient are relevant for attribution patching.
Figure 8: Pearson correlation coefficients between different attribution patching methods and the activation patching method. The values are averaged for different models (GPT-2 family), and the error bars represent the standard deviation. This metric does not account for amplitude differences between the methods.